Prove the following identities



(a)(cos t)(tan t)(csc t) =1


1
Expert's answer
2022-01-14T12:44:56-0500

It follows that

"\cos t\cdot\tan t\cdot\csc t\\
 =\cos t\cdot\frac{\sin t}{\cos t}\cdot\frac{1}{\sin t}\\=
\frac{\cos t\cdot\sin t}{\cos t\cdot\sin t}=1"